OpenStudy (anonymous):

-3(x-6)=4x+4 show work?

thomaster (thomaster):

First distribute the -3(x-6)

thomaster (thomaster):

Then add or subtract like terms on both sides. x to the left and numbers to the right. then divide the number side by the amount of x's you have

OpenStudy (anonymous):

Can you show me how to do that?

thomaster (thomaster):

 So we now have -3x+18=4x+4 now we need to get the x's to the left, so we subtract 4x from both sides. that would leave -7x+18=4 now subtract 18 on both sides to get rid of the +18 on the left. that would give -7x=-14 Now divide the number side by the amount of x's you have that would be \(\sf\dfrac{-14}{-7}=2\)